OpenStudy (anonymous):

I need to solve using the elimination method 5x + 5y = -13 7x - 3y =17

OpenStudy (anonymous):

multiply first by 7 : 35x + 35y = -91 ----(A) multiply second by 5 : 35x - 15y = 85 ----(B) (A) - (B) gives 50y = - 176 (We have eliminated x terms here) so y = -176/50 = -88/25 Put y = -88/5 in your original second equation to get 7x -3(-88/5) = -13 7x + 264/5 = -13 7x = -13 - 264/5 7x = -329/5 x = -329/35 x = -47/5 so final answer is x = -47/5 and y = -88/25

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I hope u hv understood the method The aim is to eliminate one of the variables so that only one is left once we have the value of one variable we can find the value of the other too

OpenStudy (anonymous):

pls click GOOD ANSWER button to show appreciation and to mark the question as solved and closed